# “Infinitely Often” Transcendence of Gamma-Function Derivatives

### v1

@@ 2.1. Values of $\boldsymbol{q\in\tfrac{1}{2}\mathbb{Z}\setminus\mathbb{Z}_{\leq 0}}$
The following theorem, which applies to all sequences $\left\{\Gamma^{\left(n\right)}\left(q\right)\right\}_{n\geq 1}$ for $q\in\mathbb{Q}\setminus\tfrac{1}{2}\mathbb{Z}$, is proved by arguments similar to those used in the proof of Theorem 1, but restricted by the fact that Euler’s reflection formula does not generalize to $\Gamma\left(q+t\right)\Gamma\left(q-t\right)$ in this case. For $q\in\mathbb{Q}\setminus\tfrac{1}{2}\mathbb{Z}$, the reflection formula naturally couples the points $q$ and 1q1-q, leading to a disjunctive result for the sequences $\left\{\Gamma^{\left(n\right)}\left(q\right)\right\}_{n\geq 1}$ and $\left\{\Gamma^{\left(n\right)}\left(1-q\right)\right\}_{n\geq 1}$.
@@ 3. Conclusion
The present study has pursued a focused objective: generalizing the “infinitely often” transcendence results of Powers [2] to $\Gamma^{\left(n\right)}\left(q\right)$ for values $q\in\mathbb{Q}\setminus\mathbb{Z}_{\leq 0}$ beyond $q=1$. This was accomplished for $q\in\tfrac{1}{2}\mathbb{Z}\setminus\mathbb{Z}_{\leq 0}$ in Theorem 1, and partially accomplished for $q\in\mathbb{Q}\setminus\tfrac{1}{2}\mathbb{Z}$ in Theorem 2.
Natural directions for further research include: (i) improving Theorem 2 by removing the disjunctive limitation; and (ii) ascertaining lower bounds on the asymptotic densities of the transcendental terms $\Gamma^{\left(n\right)}\left(q\right)$ for $q\in\tfrac{1}{2}\mathbb{Z}\setminus\mathbb{Z}_{\leq 0}$. The latter objective, although challenging, appears feasible because Theorem 1(iv) of Powers [2] was able to provide such a lower bound for the sequence $\left\{\Gamma^{\left(n\right)}\left(1\right)\right\}_{n\geq 1}$, conditional on a conjecture rigorously certified by computer for all $n\in\left\{2,3,\ldots,26\right\}$.

### v2

@@ 3. Values of $q\in\mathbb{Q}\setminus\tfrac{1}{2}\mathbb{Z}$
In this section, we consider transcendental elements in sequences $\left\{\Gamma^{\left(n\right)}\left(q\right)\right\}_{n\geq 1}$ for $q\in\mathbb{Q}\setminus\tfrac{1}{2}\mathbb{Z}$. Although the approach is similar to that of Section 2.1, it is restricted by the fact that Euler’s reflection formula (in conjunction with the functional equation) does not yield a simple closed form for $\Gamma\left(q+t\right)\Gamma\left(q-t\right)$ in this case. Instead, the reflection formula naturally couples the points $q$ and 1q1-q, leading to a disjunctive result for the sequences $\left\{\Gamma^{\left(n\right)}\left(q\right)\right\}_{n\geq 1}$ and $\left\{\Gamma^{\left(n\right)}\left(1-q\right)\right\}_{n\geq 1}$.
@@ 4. Conclusion
In the present study, we investigated the “infinitely often” transcendence of $\Gamma^{\left(n\right)}\left(q\right)$ for $q\in\mathbb{Q}\setminus\mathbb{Z}_{\leq 0}$. Theorem 1 showed that for all $q\in\tfrac{1}{2}\mathbb{Z}\setminus\mathbb{Z}_{\leq 0}$, the sequence $\left\{\Gamma^{\left(n\right)}\left(q\right)\right\}_{n\geq 1}$ contains infinitely many transcendental elements, and Theorem 2 gave a lower bound, $\beta\left(N\right)\asymp N^{-1/2}$, for the density of transcendental elements $\Gamma^{\left(n\right)}\left(q\right)$ among $n\in\left\{1,2,\ldots,N\right\}$. Although it was not possible to extend the full implications of Theorem 1 to $q\in\mathbb{Q}\setminus\tfrac{1}{2}\mathbb{Z}$, we were able to provide a weaker (disjunctive) result, showing that at least one of the sequences $\left\{\Gamma^{\left(n\right)}\left(q\right)\right\}_{n\geq 1}$, $\left\{\Gamma^{\left(n\right)}\left(1-q\right)\right\}_{n\geq 1}$ contains infinitely many transcendental elements.
Natural directions for further research include: (i) improving Theorem 3 by removing the disjunctive limitation; and (ii) for fixed $n\in\mathbb{Z}_{\geq 1}$ and $\kappa\in\left(0,1\right)\cap\mathbb{Q}$, ascertaining lower bounds on the densities of transcendental elements $\Gamma^{\left(n\right)}\left(m+\kappa\right)$ among $\left|m\right|\in\left\{0,1,\ldots,M\right\}$. Preliminary progress toward the latter objective is presented in Powers [14].